A card is drawn at random from a 52-card deck. What is the probability that the card is an ace or a spade?
Answer by ewatrrr(24785)


Hi
A card is drawn at random from a 52-card deck.
P( the card is an ace 'or' a spade) = P(ace) + P(spade) - P(ace of spades)
P( the card is an ace 'or' a spade) = 4/52 + 13/52 - 1/52 = 16/52 = 4/13